Last week Ava and John sold $1225 in goods if John sold three times more than Ava how much did Ava sell ?
Answer by ikleyn(52833)

.

Let x = how much did Ava sell (in dollars).


Then John sold  4x,  according to the condition.


x + 4x = 1225

5x     = 1225

 x     = 1225/5 = 245.


ANSWER.  245 dollars.

Solved.